OpenStudy (anonymous):

An infinite geometric series with second term 9/4 has a sum of 81/8. Determine the possible values for the first term. I understand the formula for an infinite geometric series is \[S=\frac{a}{1-r}\] Given that, I need to find a and r. I thought I could find r by using \[t_n=ar^{n-1}\] but I still do not know r or a. I would know how to do this if I was given two terms, but not when I am given only one.

OpenStudy (anonymous):

Well. s = 81/8 => 8a = 81 - 81r --1 And, ar = 9/4

OpenStudy (anonymous):

You have two equations. Two unkowns. What is the problem?

OpenStudy (anonymous):

i am also confused, and i think the problem is ill posed. it could be \[\sum_{k=0}^{\infty}ar^k\] or \[\sum_{k=1}^{\infty}ar^k\]

OpenStudy (anonymous):

@satellite73 - I am not sure. It does not give an equation in sigma notation. I have basically typed out the question as it is. @siddhantsharan - What are you doing on the second line? => 8a = 81 - 81r --1

OpenStudy (anonymous):

Why? From the sum, thing, we have the first equations. 8a = 81 - 81r Now the second term is First term * Common Ratio. = ar = 9/4

OpenStudy (anonymous):

The sum of the infinite series is, S = a/ (1-R) I just plugged in s.

OpenStudy (anonymous):

Oh, I guess the --1 part was a mistake then.

OpenStudy (anonymous):

Oh yeahh. I was just writing ------equation(1) And ar = 9/4 as equation 2.

OpenStudy (anonymous):

27/8 and 27/4 are the possible values for the first term?

OpenStudy (anonymous):

Wouldn't 8a = 81-81r simplify to 8a = 81(1-r), then 8a/81=(1-r), then (1-r)/a?

OpenStudy (anonymous):

It would. My original equation was : \[81/8 = a/(1-r)\]

OpenStudy (anonymous):

Wait a sec. Can you list both the equations you used again? I thought the second one was 8a = 81 - 81r and first ar=9/4.

OpenStudy (anonymous):

I plugged in ar=9/4, a=27/8 to find r = 2/3. Putting that in S=a/(1-r) gave me -27/10.

OpenStudy (anonymous):

\[Equation 1 : 81/8 = a/1-r\] \[Equation 2 = ar = 9 /4\]

OpenStudy (anonymous):

If. ar = 9/4, a = 27/8, r = 2/3 S = 27/8 / 1/3 = 81/8

OpenStudy (anonymous):

Which clearly satisfies the condition!

OpenStudy (anonymous):

Woops, plugged in the wrong number. Still don't understand how you found a though. Did you just solve \[\frac{a}{1-r}-\frac{81}{8}=ar-\frac{9}{4}\]

OpenStudy (anonymous):

No. From ar = 9/4 ..... => r = 9/4a Substituted r in the first equation. Got a quadratic in a. Solved for 2 values.

OpenStudy (anonymous):

\[Quadratic was : 32a^2 - 324a + 729 = 0\]

OpenStudy (anonymous):

Look. \[r = 9/4a\] \[8a = 81 - 81(9/4a)\] Multiply the equation by a. Take the LCM. You would get a quadratic.

OpenStudy (anonymous):

I see. Thank you very much!